# Arithmetic Properties of Several Generalized-Constant Sequences, with Implications for $\boldsymbol{\Gamma^{\left(n\right)}\left(1\right)}$

### v5

@@ Abstract.
Neither the Euler-Mascheroni constant, $\gamma=0.577215\ldots$, nor the Euler-Gompertz constant, $\delta=0.596347\ldots$, is currently known to be irrational. However, it has been proved that at least one of them is transcendental. The two constants are related through a well-known equation of Hardy, equivalent to $\gamma+\delta/e=\textrm{Ein}\left(1\right)$, which recently has been generalized to $\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e=\eta^{\left(n\right)},\>n\geq 0$ for sequences of constants $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ (given respectively by raw, conditional, and partial moments of the Gumbel(0,1) probability distribution). Investigating the $\gamma^{\left(n\right)}$ through recurrence relations (where $\gamma^{\left(0\right)}=1$ and $\gamma^{\left(1\right)}=\gamma$), we find that at least one of the pair $\{\gamma,\gamma^{\left(2\right)}\}$ and at least two of each set $\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\}$, for $2\leq n\leq 26$, are transcendental, with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$. We then show, via a theorem of Shidlovskii, that the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$, implying that at least one of each pair, $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\}$ and $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\}$, and at least two of the triple $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\}$, are transcendental for all $n\geq 1$. Further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ reveals that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often with positive asymptotic density. Finally, we provide parallel results for the sequences $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$.
@@
Euler-Mascheroni constant; Euler-Gompertz constant; Gumbel distribution; irrationality; transcendence; algebraic independence; gamma-function derivatives.
@@ 1.1. Three Fundamental Sequences
$\gamma^{\left(n\right)}+\dfrac{\delta^{\left(n\right)}}{e}=\eta^{\left(n\right)}.$ (8)
@@ 1.2. Asymptotic Behavior
$10$ $\qquad 3,627,042.4127568947\ldots\qquad$ $11.3803468877\ldots$ $\qquad 3,627,038.2261612415\ldots\qquad$
@@ 1.3. Outline of Results
In Section 2, we begin our study of the arithmetic properties of the three generalized-constant sequences by investigating recurrence relations associated with the $\gamma^{\left(n\right)}$. These recurrences reveal that at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental. As will be seen, the latter result follows from verification, for $n\in\left\{2,3,\ldots,26\right\}$, of a certain reasonable conjecture (which, if true for all $n\in\mathbb{Z}_{\geq 2}$, necessitates that the $\gamma^{\left(n\right)}$ are transcendental infinitely often). Since $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$, all conclusions for the $\gamma^{\left(n\right)}$ apply equally well to the sequence $\Gamma^{\left(n\right)}\left(1\right)$.
In Section 3, we employ a theorem of Shidlovskii (1989) to demonstrate the algebraic independence, and therefore transcendence, of the $\eta^{\left(n\right)}$ for all $n\geq 0$. This implies that at least one element of each pair, $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$. Moreover, further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ shows that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$. Finally, Section 4 addresses two additional sequences, $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$, satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$, and provides results parallel to those for the original system.
@@ 2.1. Basic Recurrences
For the case of $X\equiv X_{\textrm{G}}\sim\textrm{Gumbel}\left(0,1\right)$, the identities in (9) and (10) can be simplified by setting
@@ 2.2.1. Algebraic Structure of the $P_{n}$
The variety $\mathcal{V}_{n}$ is irreducible and has function field of transcendence degree $n+1$.
@@ 2.2.2. Existence and Non-Triviality of $\widetilde{P}_{n,\theta}=0$
$\mathbf{J}_{-\theta}^{\left(n\right)}\left(j,\ell\right)=\left[\frac{d\gamma^{\left(j\right)}}{d\kappa_{\ell}}\right],\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}$
@@
$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\left(j,\ell\right)\right|_{\mathcal{K}_{n}^{*}}=\left.\left[\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{\ell}}\right]\right|_{\mathcal{K}_{n}^{*}},$
@@
$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\left(j,\ell\right)\right|_{\mathcal{K}_{n}^{*}}=\binom{j}{\ell}$
@@
$\mathbf{M}_{-\theta}^{\left(n\right)}\left(j,\ell\right)=\binom{j}{\ell},\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}.$
To see that $\det\left(\mathbf{M}_{-\theta}^{\left(n\right)}\right)\neq 0$, we note that $\mathbf{M}_{-\theta}^{\left(n\right)}$ is a submatrix of the lower-triangular Pascal matrix,
$\mathbf{P}^{\left(n\right)}\left(r,c\right)=\binom{r}{c},\quad r\in\left\{0,1,\ldots,2n\right\},c\in\left\{0,1,\ldots,2n\right\},$
@@ 2.3. Transcendence Results
The transformation of the original polynomials of (16) into the reduced form of (17) lays the foundation for part (ii) of Theorem 1 below. However, one additional condition on the $\widetilde{P}_{n,\theta}$ is required. Although we know there exists $\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ that can be expressed as a non-trivial polynomial equation in $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$ of the form given by (19), it remains possible that $A_{d}^{\left(n,\theta\right)}=0$ for all $d\in\left\{1,2,\ldots,D_{n,\theta}\right\}$ at the actual realizations of the elements $\gamma^{\left(k\right)}\in\mathcal{G}_{n}^{\left(2n\right)}$ (to be denoted by $\boldsymbol{\gamma^{*}}\Longleftrightarrow\gamma^{\left(k\right)}=\textrm{E}_{X}\left[X_{\textrm{G}}^{k}\right]\>\forall\>k\in\mathbb{Z}_{\geq 1}$). Lemma 5 below (supported by the subsequent Conjecture 1) ensures this does not happen.
@@
$\mathbf{J}_{m,\theta}^{\left(n\right)}\left(i,j\right)=\left[\frac{\partial P_{i}}{\partial\gamma^{\left(j\right)}}\right],\quad i\in\left\{2,3,\ldots,m\right\},j\in\left(\left\{2,3,\ldots,m-1\right\}\cup\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),$
@@
$\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ (21)
@@
$\begin{array}[]{c}M_{m,\theta}^{*}\left(t\right)=u^{\widehat{q}^{\left(m,\theta\right)}}t^{D}+\beta_{D-1}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{D-1}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)t^{D-1}+\cdots\\ +\beta_{0}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{0}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)\in\mathcal{U}_{m,\theta}\left[t\right],\end{array}$
with $M_{m,\theta}^{*}\left(\theta\right)=0$ in $\left(\mathcal{S}_{m}\right)_{s}$ and leading coefficient $u^{\widehat{q}^{\left(m,\theta\right)}}$. Since this equality holds in the localization $\left(\mathcal{S}_{m}\right)_{s}$, there exists $q\in\mathbb{Z}_{\geq 0}$ such that $s^{q}M_{m,\theta}^{*}\left(\theta\right)=0$ in $\mathcal{S}_{m}$.
@@
$\left.\left(s^{q}M_{m,\theta}^{*}\left(\theta\right)\right)\right|_{\boldsymbol{\gamma^{*}}}=0\Longrightarrow\left(\left.s\right|_{\boldsymbol{\gamma^{*}}}\right)^{q}M_{m,\theta}^{*}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$
$\Longrightarrow M_{m,\theta}^{*}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0;$
and second, for the leading coefficient of $M_{m,\theta}^{*}\left(t\right)$, we find
@@
These facts imply that $M_{m,\theta}^{*}\left(t\right)$ specializes to a non-zero polynomial in $t$ that vanishes at $t=\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$, and so cannot be a constant. Thus, at least one coefficient of the polynomial, for degree $d\geq 1$, must remain non-zero at $\boldsymbol{\gamma^{*}}$.
Finally, let $\widehat{M}_{m,\theta}^{*}\left(t\right)$ be a primitive integer multiple of $M_{m,\theta}^{*}\left(t\right)$ (viewed in $\mathbb{Z}\left[\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right]\left[t\right]$), and factor $\widehat{M}_{m,\theta}^{*}\left(t\right)$ into irreducibles in this unique factorization domain. We then can choose an irreducible factor $\widetilde{P}_{m,\theta}\left(t\right)$ such that $\widetilde{P}_{m,\theta}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$ and $\left.\widetilde{P}_{m,\theta}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}\not\equiv 0$ (because $\left.\widehat{M}_{m,\theta}^{*}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}$ is a non-zero polynomial with root $\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$). This $\widetilde{P}_{m,\theta}\in\mathbb{Z}\left[\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]$ not only provides the non-trivial polynomial relation in $\theta$ required by (17), but also has the property that when written as a polynomial in $\theta$, it possesses at least one coefficient, for degree $d\geq 1$, that remains non-zero at $\boldsymbol{\gamma^{*}}$. ∎
The conjecture below, which we have verified computationally for all $n\in\left\{2,3,\ldots,26\right\}$, asserts that condition (21) of Lemma 5 is true. [^9] For each $n$, we executed a Python program (on an Apple M4 Max MacBook computer with 48 GB memory) to provide rigorous certification of the sufficient condition in (21). The program first evaluated $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ symbolically for each $m\in\left\{2,3,\ldots,n\right\}$, and then confirmed $\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ using interval arithmetic. The maximum value of $n$ tested (i.e., $26$) was simply the largest $n$ for which the program completed all checks within approximately 24 hours. Clearly, greater values of $n$ could be tested, but the program’s runtime increases approximately exponentially in $n$. (For example, $n=30$ would require a runtime of about 1 week.) In addition to its computational verification, the conjecture is realistic on theoretical grounds because it holds for any given $n$ unless $\boldsymbol{\gamma^{*}}$ is forced to lie on one of finitely many special hypersurfaces (i.e., the vanishing loci of the minors $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$).
@@
$\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$
@@
(ii) if Conjecture 1 holds for a given $n\in\mathbb{Z}_{\geq 2}$, then at least two elements of the set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental;
(iii) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\gamma^{\left(n\right)}$ is transcendental infinitely often; and

### v6

# Transcendence Results for $\boldsymbol{\Gamma^{\left(n\right)}\left(1\right)}$ and Related Generalized-Constant Sequences

@@ Abstract.
Neither the Euler-Mascheroni constant, $\gamma=0.577215\ldots$, nor the Euler-Gompertz constant, $\delta=0.596347\ldots$, is currently known to be irrational. However, it has been proved that at least one of them is transcendental. The two constants are related through a well-known equation of Hardy, equivalent to $\gamma+\delta/e=\textrm{Ein}\left(1\right)$, which recently has been generalized to $\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e=\eta^{\left(n\right)},\>n\geq 0$ for sequences of constants $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ (given respectively by raw, conditional, and partial moments of the Gumbel(0,1) probability distribution). Investigating $\gamma^{\left(n\right)}=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right),\>n\geq 1$ through recurrence relations, we find that the $\gamma^{\left(n\right)}$ are transcendental infinitely often and that at least two of each set $\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\},\>2\leq n\leq 26$ are transcendental (with analogous results for $\Gamma^{\left(n\right)}\left(1\right)$). We then show, via a theorem of Shidlovskii, that the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$, implying that at least one of each pair, $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\}$ and $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\}$, and at least two of the triple $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\}$, are transcendental for all $n\geq 1$. Further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ reveals that both the $\delta^{\left(n\right)}/e$ and $\delta^{\left(n\right)}$ are transcendental infinitely often with asymptotic densities of at least $1/2$. Finally, we provide parallel results for the sequences $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$.
@@
Euler-Mascheroni constant; Euler-Gompertz constant; Gumbel distribution; transcendence; algebraic independence; gamma-function derivatives.
@@ 1.1. Three Fundamental Sequences
$\gamma^{\left(n\right)}+\dfrac{\delta^{\left(n\right)}}{e}=\eta^{\left(n\right)},$ (8)
and note that
$\gamma^{\left(n\right)}=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right).$
@@ 1.2. Asymptotic Behavior
$10$ $3,627,042.4127568947\ldots$ $11.3803468877\ldots$ $3,627,038.2261612415\ldots$
@@ 1.3. Outline of Results
In Section 2, we begin our study of the arithmetic properties of the three generalized-constant sequences by investigating recurrence relations associated with $\gamma^{\left(n\right)}=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right)$. These recurrences reveal that the $\gamma^{\left(n\right)}$ are transcendental infinitely often and that at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental (with analogous implications for $\Gamma^{\left(n\right)}\left(1\right)$). As will be seen, the latter result follows from verification, for $n\in\left\{2,3,\ldots,26\right\}$, of a certain reasonable conjecture that, if true for all $n\in\mathbb{Z}_{\geq 2}$, enables one to derive an asymptotic lower bound on the density of transcendental terms.
In Section 3, we employ a theorem of Shidlovskii (1989) to demonstrate the algebraic independence, and therefore transcendence, of the $\eta^{\left(n\right)}$ for all $n\geq 0$. This implies that at least one element of each pair, $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$. Moreover, further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ shows that both the $\delta^{\left(n\right)}/e$ and $\delta^{\left(n\right)}$ are transcendental infinitely often, with the densities of both sets of transcendental terms having asymptotic lower bound $1/2$. Finally, Section 4 addresses two additional sequences, $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$, satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$, and provides results parallel to those for the original system.
@@ 2.1. Basic Recurrences
For the case of $X\sim F_{X}\equiv\textrm{Gumbel}\left(0,1\right)$, the identities in (9) and (10) can be simplified by setting
@@ 2.2.1. Algebraic Structure of the $P_{n}$
The variety $\mathcal{V}_{n}$ is irreducible and has a function field of transcendence degree $n+1$.
@@ 2.2.2. Existence and Non-Triviality of $\widetilde{P}_{n,\theta}=0$
$\mathbf{J}_{-\theta}^{\left(n\right)}=\left[\frac{d\gamma^{\left(j\right)}}{d\kappa_{\ell}}\right],\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}$
@@
$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\right|_{\mathcal{K}_{n}^{*}}=\left.\left[\frac{\partial\gamma^{\left(j\right)}}{\partial\kappa_{\ell}}\right]\right|_{\mathcal{K}_{n}^{*}},$
@@
$\left.\mathbf{J}_{-\theta}^{\left(n\right)}\right|_{\mathcal{K}_{n}^{*}}=\left[\binom{j}{\ell}\right$]
@@
$\mathbf{M}_{-\theta}^{\left(n\right)}=\left[\binom{j}{\ell}\right],\quad j\in\left(\left\{1,n,n+1,\ldots,2n\right\}\setminus\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),\ell\in\left\{1,2,3,5,\ldots,2n-1\right\}.$
To see that $\det\left(\mathbf{M}_{-\theta}^{\left(n\right)}\right)\neq 0$, we note that $\mathbf{M}_{-\theta}^{\left(n\right)}$ is a submatrix of the $\left(2n+1\right)\times\left(2n+1\right)$ lower-triangular Pascal matrix,
$\mathbf{P}^{\left(n\right)}=\left[\binom{r}{c}\right],\quad r\in\left\{0,1,\ldots,2n\right\},c\in\left\{0,1,\ldots,2n\right\},$
@@ 2.3. Transcendence Results
The transformation of the original polynomials of (16) into the reduced form of (17) lays the foundation for parts (iii) and (iv) of Theorem 1 below. However, one additional condition on the $\widetilde{P}_{n,\theta}$ is required. Although we know there exists $\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ that can be expressed as a non-trivial polynomial equation in $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$ of the form given by (19), it remains possible that $A_{d}^{\left(n,\theta\right)}=0$ for all $d\in\left\{1,2,\ldots,D_{n,\theta}\right\}$ at the actual realizations of the elements $\gamma^{\left(k\right)}\in\mathcal{G}_{n}^{\left(2n\right)}$ (to be denoted by $\boldsymbol{\gamma^{*}}\Longleftrightarrow\gamma^{\left(k\right)}=\textrm{E}_{X}\left[X_{\textrm{G}}^{k}\right]\>\forall\>k\in\mathbb{Z}_{\geq 1}$). Lemma 5 below (supported by the subsequent Conjecture 1) ensures this does not happen.
@@
$\mathbf{J}_{m,\theta}^{\left(n\right)}=\left[\frac{\partial P_{i}}{\partial\gamma^{\left(j\right)}}\right],\quad i\in\left\{2,3,\ldots,m\right\},j\in\left(\left\{2,3,\ldots,m-1\right\}\cup\left\{k:\theta=\gamma^{\left(k\right)}\right\}\right),$
@@
$\det\left(\left.\mathbf{J}_{m,\theta}^{\left(n\right)}\right|_{\gamma^{*}}\right)\neq 0$ (21)
@@
$\begin{array}[]{c}\widehat{M}_{m,\theta}\left(t\right)=u^{\widehat{q}^{\left(m,\theta\right)}}t^{D}+\beta_{D-1}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{D-1}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)t^{D-1}+\cdots\\ +\beta_{0}^{\left(m,\theta\right)}u^{\left(\widehat{q}^{\left(m,\theta\right)}-q_{0}^{\left(m,\theta\right)}\right)}\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right)\in\mathcal{U}_{m,\theta}\left[t\right],\end{array}$
with $\widehat{M}_{m,\theta}\left(\theta\right)=0$ in $\left(\mathcal{S}_{m}\right)_{s}$ and leading coefficient $u^{\widehat{q}^{\left(m,\theta\right)}}$. Since this equality holds in the localization $\left(\mathcal{S}_{m}\right)_{s}$, there exists $q\in\mathbb{Z}_{\geq 0}$ such that $s^{q}\widehat{M}_{m,\theta}\left(\theta\right)=0$ in $\mathcal{S}_{m}$.
@@
$\left.\left(s^{q}\widehat{M}_{m,\theta}\left(\theta\right)\right)\right|_{\boldsymbol{\gamma^{*}}}=0\Longrightarrow\left(\left.s\right|_{\boldsymbol{\gamma^{*}}}\right)^{q}\widehat{M}_{m,\theta}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$
$\Longrightarrow\widehat{M}_{m,\theta}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0;$
and second, for the leading coefficient of $\widehat{M}_{m,\theta}\left(t\right)$, we find
@@
These facts imply that $\widehat{M}_{m,\theta}\left(t\right)$ specializes to a non-zero polynomial in $t$ that vanishes at $t=\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$, and so cannot be a constant. Thus, at least one coefficient of the polynomial, for degree $d\geq 1$, must remain non-zero at $\boldsymbol{\gamma^{*}}$.
Finally, let $\widehat{M}_{m,\theta}^{\prime}\left(t\right)$ be a primitive integer multiple of $\widehat{M}_{m,\theta}\left(t\right)$ (viewed in $\mathbb{Z}\left[\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\theta\right\}\right]\left[t\right]$), and factor $\widehat{M}_{m,\theta}^{\prime}\left(t\right)$ into irreducibles in this unique factorization domain. We then can choose an irreducible factor $\widetilde{P}_{m,\theta}\left(t\right)$ such that $\widetilde{P}_{m,\theta}\left(\left.\theta\right|_{\boldsymbol{\gamma^{*}}}\right)=0$ and $\left.\widetilde{P}_{m,\theta}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}\not\equiv 0$ (because $\left.\widehat{M}_{m,\theta}^{\prime}\left(t\right)\right|_{\boldsymbol{\gamma^{*}}}$ is a non-zero polynomial with root $\left.\theta\right|_{\boldsymbol{\gamma^{*}}}$). This $\widetilde{P}_{m,\theta}\in\mathbb{Z}\left[\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right]$ not only provides the non-trivial polynomial relation in $\theta$ required by (17), but also has the property that when written as a polynomial in $\theta$, it possesses at least one coefficient, for degree $d\geq 1$, that remains non-zero at $\boldsymbol{\gamma^{*}}$. ∎
The conjecture below, which we have verified computationally for all $n\in\left\{2,3,\ldots,26\right\}$, asserts that condition (21) of Lemma 5 is true. [^9] For each $n$, we executed a Python program (on an Apple M4 Max MacBook computer with 48 GB memory) to provide rigorous certification of the sufficient condition in (21). The program first evaluated $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ symbolically for each $m\in\left\{2,3,\ldots,n\right\}$, and then confirmed $\det\left(\left.\mathbf{J}_{m,\theta}^{\left(n\right)}\right|_{\gamma^{*}}\right)\neq 0$ using interval arithmetic. The maximum value of $n$ tested (i.e., $26$) was simply the largest $n$ for which the program completed all checks within approximately 24 hours. Clearly, greater values of $n$ could be tested, but the program’s runtime increases approximately exponentially in $n$. (For example, $n=30$ would require a runtime of about 1 week.) In addition to its computational verification, the conjecture is realistic on theoretical grounds because it holds for any given $n$ unless $\boldsymbol{\gamma^{*}}$ is forced to lie on one of finitely many special hypersurfaces (i.e., the vanishing loci of the minors $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$).
@@
$\det\left(\left.\mathbf{J}_{m,\theta}^{\left(n\right)}\right|_{\gamma^{*}}\right)\neq 0$
@@
(ii) $\gamma^{\left(n\right)}$ is transcendental infinitely often;
(iii) if Conjecture 1 holds for a given $n\in\mathbb{Z}_{\geq 2}$, then at least two elements of the set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental; and